# Graph Neural Networks for Next-Generation-IoT: Recent Advances and Open Challenges

### v1

@@ Abstract
Graph Neural Networks (GNNs) have emerged as a critical tool for optimizing and managing the complexities of the Internet of Things (IoT) in next-generation networks. This survey presents a comprehensive exploration of how GNNs may be harnessed in 6G IoT environments, focusing on key challenges and opportunities through a series of open questions. We commence with an exploration of GNN paradigms and the roles of node, edge, and graph-level tasks in solving wireless networking problems and highlight GNNs’ ability to overcome the limitations of traditional optimization methods. This guidance enhances problem-solving efficiency across various next-generation (NG) IoT scenarios. Next, we provide a detailed discussion of the application of GNN in advanced NG enabling technologies, including massive MIMO, reconfigurable intelligent surfaces, satellites, THz, mobile edge computing (MEC), and ultra-reliable low latency communication (URLLC). We then delve into the challenges posed by adversarial attacks, offering insights into defense mechanisms to secure GNN-based NG-IoT networks. Next, we examine how GNNs can be integrated with future technologies like integrated sensing and communication (ISAC), satellite-air-ground-sea integrated networks (SAGSIN), and quantum computing. Our findings highlight the transformative potential of GNNs in improving efficiency, scalability, and security within NG-IoT systems, paving the way for future advances. Finally, we propose a set of design guidelines to facilitate the development of efficient, scalable, and secure GNN models tailored for NG IoT applications.
@@ I Introduction
The Internet of Things (IoT) has revolutionized the way we interact with our environment, supporting a vast network of interconnected devices that communicate and exchange data seamlessly [1, 2]. As we move towards next-generation networks, the integration of IoT within this advanced framework, termed as NG-IoT, promises unprecedented improvements in connectivity, latency, and data throughput [3]. All the advantages accrue from sophisticated technologies, including massive multiple-input multiple-output (MIMO) schemes, reconfigurable intelligent surfaces (RIS), satellites, terahertz (THz), mobile edge computing (MEC), ultra-reliable low-latency communications (URLLC), and blockchain. Despite their potential, this evolution presents several open challenges, including the need for efficient resource allocation, robust data handling, and effective model-driven approaches. Future technologies such as integrated sensing and communication (ISAC), satellite-air-ground-sea integrated networks (SAGSINs), and quantum computing will further expand the realm of possibilities in NG-IoT networks.
@@
Optimization and data/model-driven approaches have been at the forefront of efforts to enhance 6G-IoT technologies [13, 14, 15, 16, 17, 18]. While these approaches have yielded notable improvements, these are often attained at the cost of evaded scalability and excessive complexity. For instance, MEC and massive MIMO require sophisticated resource allocation and interference management, which are computationally intensive and, hence, challenging to implement in real-time scenarios. Similarly, the deployment of RIS and THz communications necessitates precise environmental adaptation and robust signal processing techniques, which traditional optimization methods struggle to handle efficiently.
Graph Neural Networks (GNNs) have emerged as a promising technique for circumventing these limitations, offering a novel approach to modeling the complex relationships and dependencies inherent in NG-IoT networks. In contrast to traditional methods, GNNs leverage graph-structured data to capture the intricate connectivity patterns of network elements. This capability allows GNNs to efficiently model the non-Euclidean data structures typical in wireless networks. GNNs can handle varying-sized inputs, making them ideal for NG-IoT networks where the number of devices and connections can change frequently [19, 20]. While traditional models usually struggle in dynamic network environments, GNNs operate on graph structures that are capable of naturally adapting to fluctuations in network topology. Hence, GNNs are capable of integrating new devices or changes in network structure without needing retraining while maintaining consistent performance even as the network evolves. To better understand the evolution of GNNs and their capabilities, Table I provides an overview of the development from early graph theory to the advanced GNN models.
GNNs also feature decentralized implementation, where computations are distributed across multiple nodes in the network rather than being centralized. This decentralization reduces computational complexity and enhances robustness against time-variant network conditions. Each node can perform computations locally and only share the absolutely necessary information with neighboring nodes, hence reducing the overall computational load and making the system more resilient to outages in the face of network structure uncertainties. By contrast, traditional deep learning methods often require extensive labeled data and centralized processing, leading to excessive computational costs and reduced flexibility. As a result, GNNs support efficient and scalable processing, making them well-suited for the evolving demands of NG-IoT networks. To illustrate the comparative advantages of GNNs, we present a comparison table (Table II) that contrasts the efficiency of optimization approaches, of traditional deep learning technologies, and of GNNs concerning their key parameters such as scalability, computational complexity, and practical implementation. This comparison highlights the superior performance of GNNs in addressing the challenging requirements of NG-IoT networks.
@@
Existing surveys have significantly contributed to the understanding of GNNs and their application in wireless communications and the IoT. For example, Lee et al. [21] discusses the potential of GNNs in wireless communications, focusing on how graphical models are constructed and their application in wireless networks. The paper Ivanov et al. [22] provides insights into resource allocation using GNNs for integrated space and terrestrial networks, while the author in Tam et al. [23] provides a review of GNN applications in areas such as network management, offloading strategies, routing optimization, virtual network function orchestration, and resource allocation. Additionally, Suarez et al. [24] focuses on specific use cases of GNNs in communication networks. The most recent survey was conducted in [25], where Sabarish et al. have explored GNN applications in IoT networks, highlighting advances in spectrum awareness, data fusion, and network intrusion detection. While these surveys have made considerable progress, this treatise provides an up-to-date critical appraisal of the relevant follow-up advances.
Although some offer insights into the application of GNNs, they often do not provide sufficient practical guidance on how and when to apply GNN techniques to specific wireless networking tasks. Additionally, wireless networks, due to their open communication nature, are particularly vulnerable to threats like eavesdropping and jamming attacks [26, 27, 28]. Nevertheless, security concerns, especially adversarial attacks, are often inadequately addressed, leaving a critical gap in understanding the robustness of GNNs in hostile environments. Moreover, prior review papers tend to focus on specific aspects of GNN applications, missing a comprehensive perspective that integrates multiple technologies and their interactions. The future potential and emerging applications of GNNs are frequently overlooked, limiting the scope to existing technologies. By contrast, to fully harness the potential of GNNs in NG-IoT networks, this paper provides an integrated survey addressing the full range of challenges and opportunities that GNNs present for NG technologies. In this paper, we present eight open questions that address key challenges and opportunities for GNNs in NG-IoT networks. These questions cover fundamental GNN paradigms, their diverse applications in NG environments, the challenges of adversarial attacks and corresponding defenses, as well as the integration of GNNs with emerging technologies such as future integrated networks and quantum computing. Each question is explored in depth, providing an overview of the existing research, challenges, and future directions. This comprehensive approach aims for highlighting the various ways GNNs can support and enhance the development of NG wireless systems. A summary of these open questions is provided in Table III. The main contributions of this paper are summarized as follows:
@@
1) How do GNN paradigms represent and process graph-structured data across different types of graphs? Graph Neural Networks (GNNs) include several variants, such as GCN, GAT, and GAEs Graph data, GNN Paradigm, and GNN variants [29, 30, 31] The explainability and scalability of GNNs [32, 33] Model depth and over-smoothing Scalability of GNN models Security
2) How do different task levels (node, edge, graph) contribute to solving NG-IoT network problems? GNN tasks can operate at node, edge, and graph levels, providing different perspectives for solving complex network challenges, such as resource allocation and link prediction in NG Node-level task [34, 35, 36, 37, 38, 39, 40, 41, 42] Edge-level task [43, 44, 45, 46, 47] Graph-level task [48, 49, 50, 51] Scalability High computational and memory requirements. Divide large graphs into subgraphs for local learning and then aggregate
@@
3) How do GNNs enhance communication and computation efficiency in massive MIMO, RIS systems, satellites, THz, MEC, and URLLC systems? Massive MIMO, RIS Systems, Satellite, THz Communication, MEC, and URLLC are essential for enhancing NG-IoT system performance, coverage, and efficiency Massive MIMO [20, 52, 53, 54, 45, 55, 56, 57] RIS Systems [58, 59, 60, 57] Satellite [61, 62, 63, 64, 65, 66, 62] THz Communication [67, 68, 69] MEC [70, 71, 72, 73, 74, 75, 76, 77, 78] URLLC [79, 80, 81, 82] Heterogeneity of network’s entities Dynamic resource allocation with diverse constraints Scalability
@@
4) How can GNNs enhance integrity, security, and scalability along with blockchain for IoT systems? Blockchain is a promising technology for ensuring data integrity and security in IoT systems. GNNs can help enhance blockchain-based IoT systems by improving node classification, enhancing security, and supporting scalable solutions User privacy [83] Application distribution among IoT networks [84] Malicious node detection [85] Computational and communication burden Ensuring data integrity and security in decentralized environments Lightweight GNN models design Privacy-preserving mechanisms
@@
5) How do adversarial attacks exploit vulnerabilities in GNN-based NG-IoT networks? GNN models are vulnerable to adversarial attacks, where small changes to input data can degrade performance. In NG-IoT networks, these attacks threaten data integrity, disrupt services, and pose security risks in various applications. Adversarial Homogeneous Graph Neural Network [86, 87, 88, 89, 90] Adversarial Heterogeneous Graph Neural Network [91, 92] Adversarial HoGNN defense Trade-off between deconstructing performance and computational complexity Explore joint attack methods involving multiple algorithms.
@@
6) What are the most effective defense techniques against adversarial attacks in GNN-based NG-IoT networks? Defense strategies are crucial to safeguarding GNNs from adversarial threats to ensure robust deployment of GNNs in critical NG-IoT applications, including smart cities, healthcare, and autonomous systems Adversarial Homogeneous GNN defense [93, 94] Adversarial Heterogeneous GNN defense [95, 96] Adversarial training Defensive distillation: Distill knowledge from a complex model to a simple one Hybrid defense approaches to improve robustness.
@@
7) How can GNNs enhance the performance and scalability of future integrated sensing and communication and space-air-ground-sea integrated networks? Future integrated networks like SAGSINs and ISAC are crucial for ensuring seamless connectivity across multiple domains. GNNs can enhance the performance and scalability of these networks by optimizing communication and sensing processes and handling complex cross-domain interactions. Integrated Communications and Sensing [97, 69] Space-air-ground-sea integrated networks [75] Scalability and heterogeneity in ISAC and SAGSINs Privacy and security issues in integrated communication and sensing Lack of standardized protocols for seamless integration of GNNs
@@
8) How can GNNs and future computational technologies, like quantum computing, work together to enhance the capabilities of NG-IoT networks? Quantum computing has the potential to address the computational limitations of GNNs, providing enhanced capabilities for processing complex graph data in NG-IoT networks Quantum computing and quantum circuits [98, 99, 100] Variational quantum circuit for GNN [101] Hybrid quantum graph neural network [102] Noisy intermediate-scale quantum (NISQ) devices Standard optimal variational quantum circuits for implementing QGNNs guaranteeing efficient circuit design and scalability Quantum cryptography with GNN
@@
  • We provide an overview of graph neural networks in NG IoT systems, discussing their definitions, paradigms, and key use cases. We explain the roles and benefits of node-, edge-, and graph-level tasks in wireless network problems, offering practical examples to guide their application in various NG-IoT scenarios.
  • We compare the performance of different GNN models, including our proposed hybrid quantum GNN harnessed for power allocation in cell-free massive MIMO, demonstrating the efficiency of GNNs and the potential of quantum GNNs for future research.
  • We present a comprehensive review of GNN applications, carefully categorized by the core technologies driving NG advances. These include massive MIMO schemes, RIS, Satellite, THz communications, MEC, URLLC, and blockchain.
@@
  • Table I offers a detailed comparison between our work and other state-of-the-art surveys in the field, emphasizing the unique contributions and advances presented in this paper.
@@
Paper Organization: The rest of the paper is organized as follows. In Section II, we provide an overview of GNNs, covering their definitions, paradigms, and the different levels of graph tasks—node, edge, sub-graph, and graph-level tasks. We discuss how these tasks can be applied to address wireless network problems and include a simulation of power allocation in cell-free massive MIMO systems to compare the performance of various GNN models. Section III presents a detailed survey of GNN applications across key NG technologies, including massive MIMO scheme, RIS, satellite communication, THz, MEC, URLLC, and blockchain. Section IV delves into adversarial GNNs, examining potential attack methods and defense strategies in GNN-based NG networks, with a focus on both homogeneous and heterogeneous graph settings. In Section V, we discuss the integration of GNNs with emerging NG technologies, including ISAC, SAGSINs, and quantum graph neural networks. Section VI will provide a set of generic design guidelines. Finally, Section VII concludes the paper. For convenience, we treat each of the eight open questions in the form of an identical structure: 1) Background; 2) State-of-the-art; 3) Challenges and future directions.
II Graph Neural Network

### v2

@@ Abstract
Graph Neural Networks (GNNs) have emerged as a powerful framework for modeling complex interconnected systems, hence making them particularly well-suited to address the growing challenges of next-generation Internet of Things (NG-IoT) networks. Despite increasing interest in this area, existing studies remain fragmented, and there is a lack of comprehensive guidance on how GNNs can be systematically applied to NG-IoT systems. As NG-IoT systems evolve toward 6G, they incorporate diverse technologies such as massive MIMO, reconfigurable intelligent surfaces (RIS), terahertz (THz) communication, satellite systems, mobile edge computing (MEC), and ultra-reliable low-latency communication (URLLC). These advances promise unprecedented connectivity, sensing, and automation but also introduce significant complexity, requiring new approaches for scalable learning, dynamic optimization, and secure, decentralized decision-making. This survey provides a comprehensive and forward-looking exploration of how GNNs can empower NG-IoT environments structured as ten open research questions that span the relevant theoretical foundations, practical deployments, and emerging integration pathways. We commence by exploring the fundamental paradigms of GNNs and articulating the motivation for their use in NG-IoT networks. Besides, to further justify their suitability, we intrinsically connect GNNs for the first time with the family of low-density parity-check (LDPC) codes, modeling the NG-IoT as dynamic constrainted graphs where GNNs harness belief propagation for convergence and interpretability through density evolution and EXIT charts. We highlight the distinct roles of node-, edge-, and graph-level tasks in tackling key challenges and demonstrate the GNNs’ ability to overcome the limitations of traditional optimization methods. Following this, we examine the application of GNNs across core NG-enabling technologies and their integration with distributed frameworks to support privacy-preservation and distributed intelligence. We then delve into the challenges posed by adversarial attacks, offering insights into defense mechanisms to secure GNN-based NG-IoT networks. Lastly, we examine how GNNs can be integrated with emerging technologies like integrated sensing and communication (ISAC), satellite-air-ground-sea integrated networks (SAGSIN), and quantum computing. Our findings highlight the transformative potential of GNNs in improving efficiency, scalability, and security within NG-IoT systems, paving the way for future advances. Finally, we summarize the key lessons learned throughout the paper and outline promissing future research directions, along with a set of design guidelines aimed at facilitating the development of efficient, scalable, and secure GNN models tailored for NG-IoT applications.
@@ I Introduction
The Internet of Things (IoT) has revolutionized the way we interact with our environment, supporting a vast network of interconnected devices that communicate and exchange data seamlessly [1, 2]. As we move toward the next-generation IoT (NG-IoT) paradigm, driven by the emergence of 6G, the integration of technologies such as massive MIMO, reconfigurable intelligent surfaces (RIS), satellites, terahertz (THz) communications, mobile edge computing (MEC), and ultra-reliable low-latency communication (URLLC) opens up unprecedented possibilities for intelligent, low-latency, and context-aware connectivity [3]. At the same time, recent advances like integrated sensing and communication (ISAC), satellite-air-ground-sea integrated networks (SAGSINs), and quantum-enhanced computing will significantly expand the system scale, heterogeneity, and complexity of NG-IoT environments. Despite their potential, the rapid development of NG-IoT networks involves higher complexity and dynamism, which introduce critical challenges in terms of scalability, robustness, and real-time decision-making. As network topologies become increasingly heterogeneous and time-varying, achieving efficient resource allocation, topology-aware learning, and resilient inference under dynamic wireless environments has become essential.
@@
Traditional approaches, such as optimization-based methods and conventional deep learning, have been at the forefront of efforts to enhance 6G-IoT technologies [13, 14, 15, 16, 17, 18]. However, these methods are increasingly inadequate in addressing these rapidly evolving multifaceted demands. While these approaches have yielded notable improvements, these are often attained at the cost of eroded scalability and excessive complexity. For instance, MEC and massive MIMO require sophisticated resource allocation and interference management, which are computationally intensive and, hence, challenging to implement in real-time scenarios. Similarly, the deployment of RIS and THz communications necessitates agile environmental adaptation and robust signal processing techniques, which traditional optimization methods struggle to handle efficiently.
Graph Neural Networks (GNNs) have emerged as a promising technique of circumventing these limitations, offering a novel approach to modeling the complex relationships and dependencies inherent in NG-IoT networks. Unlike traditional optimization or deep learning methods that operate on structured, Euclidean data, GNNs are designed to learn over irregular, non-Euclidean domains such as graphs. This allows GNNs to naturally represent the complex, multi-hop relationships among entities in wireless networks, such as IoT devices, edge servers, UAVs, and base stations, where the data structure is inherently graph-like [19, 20]. To better understand the evolution of GNNs and their capabilities, Table I highlights their evolution from early graph theory to advanced GNN models.
Furthermore, GNNs can dynamically adapt to evolving topologies without retraining, handle varying-sized networks, and support decentralized inference, all of which are critical for NG-IoT deployments. Their message-passing architecture captures both local and global contexts, enabling robust predictions even under partial observability or uncertain channel conditions. By contrast, traditional optimization techniques often suffer from high complexity and limited generalization capability, while deep neural networks lack the inductive relational biases required for graph-structured environments. To illustrate this contrast, Table II provides a comparative analysis of optimization-based approaches, conventional deep learning, and GNNs in terms of scalability, computational complexity, integration flexibility, and real-world applicability.
@@
Existing surveys have significantly contributed to the understanding of GNNs and their application in wireless communications and in the IoT. For example, Lee et al. [21] discussed the potential of GNNs in wireless communications, focusing on how graphical models are constructed and on their application in wireless networks. As a further advance, Ivanov et al. [22] provided insights into resource allocation using GNNs for integrated space and terrestrial networks, while Tam et al. [23] provided a review of GNN applications in areas such as network management, offloading strategies, routing optimization, virtual network function orchestration, and resource allocation. Additionally, Suarez et al. [24] focused on specific use cases of GNNs in communication networks. The most recent survey was offered by Sabarish et al. [25], who have explored GNN applications in IoT networks, highlighting their benefits in spectrum awareness, data fusion, and network intrusion detection. While these surveys have made considerable progress, this treatise provides an up-to-date critical appraisal of the relevant follow-up advances. Besides, these surveys tend to focus on isolated use cases or specific layers of the network stack, lacking a holistic exploration of how GNNs can address the growing complexity of next-generation IoT (NG-IoT) systems. Most prior studies provided technical summaries, but fall short of offering a practical framework or decision-making guide for selecting appropriate GNN models for different NG-IoT tasks.
Moreover, critical issues such as deployment feasibility, system cost, energy efficiency, and communication overhead, which are essential for real-world GNN deployment, have not been discussed. The role of decentralized structures and federated learning, which are increasingly important in privacy-sensitive distributed IoT networks, also remains underexplored. Additionally, wireless networks, due to their open and distributed nature, are particularly vulnerable to threats such as eavesdropping, jamming, and adversarial attacks [26, 27, 28]. Yet, most surveys inadequately tackle GNN robustness under such threats, leaving a critical knowledge gap in the context of secure and resilient NG-IoT operations.
Furthermore, prior review papers tend to focus on specific aspects of GNN applications, missing a holistic perspective that integrates multiple technologies and their interactions. The future potential and emerging applications of GNNs are frequently overlooked, limiting the scope to existing technologies. Importantly, the question of why GNNs are needed in NG-IoT networks has often been overlooked in previous studies. In contrast, this paper aims to comprehensively address this gap by presenting an integrated survey that explores both the theoretical and practical dimensions of GNN deployment in next-generation IoT systems. We formulate ten open questions that collectively examine the core challenges and opportunities of applying GNNs in NG-IoT environments. These include foundational discussions, including the justification for GNNs based on the structural characteristics of NG-IoT networks and a novel theoretical bridge to low-density parity-check (LDPC) codes, as well as strategies for model selection, convergence analysis, and interpretability. Moreover, we investigate real-world deployment challenges, including scalability, cost-efficiency, security, and application diversity, along with the role of GNNs in enabling distributed and federated learning across heterogeneous network infrastructures. We also examine how GNNs can be integrated with future-proof technologies such as integrated sensing and communication, space-air-ground-sea integrated networks, and quantum computing. Each question is explored in depth, providing an overview of the existing body of research, open challenges, and future directions. Through this lens, we aim to equip researchers and practitioners with a holistic understanding of how GNNs can empower the next wave of intelligent IoT systems. A summary of these open questions is provided in Table III. The main contributions of this paper are summarized as follows:
@@
1) What Are Graphs and Graph Neural Networks? Graph definition and Graph Neural Networks (GNNs) paradism Graph data, GNN Paradigm, and GNN variants [29, 30, 31] The explainability and scalability of GNNs [32, 33] Model depth and over-smoothing Scalability of GNN models Security
2) Why GNNs for NG-IoT Networks? How Do We Model NG-IoT Systems as Graphs? NG-IoT networks feature dynamic topology, heterogeneity, and non-Euclidean structures, making them well-suited for graph modeling. GNNs enable effective learning at node, edge, and graph levels to address these complexities. Node-level task [34, 35, 36, 37, 38, 39, 40, 41, 42] Edge-level task [43, 44, 45, 46, 47] Graph-level task [48, 49, 50, 51] High computational and memory requirements. Divide large graphs into subgraphs for local learning and then aggregate Adapting DE and EXIT chart for analyze GNN and model design
3) How Can GNNs Be Efficiently Implemented and Optimized for Real-World NG-IoT Systems? Deploying GNN in real NG-IoT environments is affected by multiple aspects, including model deployment, multi-modal data, socio-economic dimentions of GNN including cost consideration and sustainability. Model Optimization for Edge and Cloud Deployment [52, 53, 54, 55] Multi-modal data processing in NG-IoT networks [56, 57] Cost consideration [58, 59, 60, 61] Resource constraints at edge devices Need for real-time, scalable inference Multimodal data integration Deployment cost and energy efficiency
@@
4) How do GNNs enhance communication and computation efficiency in massive MIMO, RIS systems, satellites, THz, MEC, and URLLC systems? Massive MIMO, RIS Systems, Satellite, THz Communication, MEC, and URLLC are essential for enhancing NG-IoT system performance, coverage, and efficiency Massive MIMO [20, 62, 63, 64, 45, 65, 66, 67] RIS Systems [68, 69, 70, 67] Satellite [71, 72, 73, 74, 75, 76, 72] THz Communication [77, 78, 79] MEC [80, 81, 82, 83, 84, 85, 86, 87, 88] URLLC [89, 90, 91, 92] Heterogeneity of network’s entities Dynamic resource allocation with diverse constraints Scalability
@@
5) How can GNNs enhance integrity, security, and scalability along with blockchain for IoT systems? Blockchain is a promising technology for ensuring data integrity and security in IoT systems. GNNs can help enhance blockchain-based IoT systems by improving node classification, enhancing security, and supporting scalable solutions User privacy [93] Application distribution among IoT networks [94] Malicious node detection [95] Computational and communication burden Ensuring data integrity and security in decentralized environments Lightweight GNN models design Privacy-preserving mechanisms
6) How Can GNNs Collaborate with Distributed Systems in NG-IoT Networks? Distributed learning and GNNs together enable scalable, privacy-aware intelligence across NG-IoT edge devices. Distributed architecture for GNN in NG-IoT networks [96, 97, 98, 99, 100, 101, 102, 103, 104] GNN for the distributed systems in NG-IoT networks [105, 38, 106, 92] Unstable training and poor convergence. Lack scalability in large or dense networks. Synchronization and consistency Evolving graphs demand online updates due to varying environment.
@@
7) How Do Adversarial Attacks Exploit Vulnerabilities in GNN-Based NG-IoT Networks? GNN models are vulnerable to adversarial attacks, where small changes to input data can degrade performance. In NG-IoT networks, these attacks threaten data integrity, disrupt services, and pose security risks in various applications. Adversarial Homogeneous Graph Neural Network [107, 108, 109, 110, 111] Adversarial Heterogeneous Graph Neural Network [112, 113] Adversarial HoGNN defense Trade-off between deconstructing performance and computational complexity Explore joint attack methods involving multiple algorithms.
@@
8) What are the most effective defense techniques against adversarial attacks in GNN-based NG-IoT networks? Defense strategies are crucial to safeguarding GNNs from adversarial threats to ensure robust deployment of GNNs in critical NG-IoT applications, including smart cities, healthcare, and autonomous systems Adversarial Homogeneous GNN defense [114, 115] Adversarial Heterogeneous GNN defense [116, 117] Adversarial training Defensive distillation: Distill knowledge from a complex model to a simple one Hybrid defense approaches to improve robustness.
@@
9) How can GNNs enhance the performance and scalability of future integrated sensing and communication and space-air-ground-sea integrated networks? Future integrated networks like SAGSINs and ISAC are crucial for ensuring seamless connectivity across multiple domains. GNNs can enhance the performance and scalability of these networks by optimizing communication and sensing processes and handling complex cross-domain interactions. Integrated Communications and Sensing [118, 79] Space-air-ground-sea integrated networks [85] Scalability and heterogeneity in ISAC and SAGSINs Privacy and security issues in integrated communication and sensing Lack of standardized protocols for seamless integration of GNNs
@@
10) How can GNNs and future computational technologies, like quantum computing, work together to enhance the capabilities of NG-IoT networks? Quantum computing has the potential to address the computational limitations of GNNs, providing enhanced capabilities for processing complex graph data in NG-IoT networks Quantum computing and quantum circuits [119, 120, 121] Variational quantum circuit for GNN [122] Hybrid quantum graph neural network [123] Noisy intermediate-scale quantum (NISQ) devices Standard optimal variational quantum circuits for implementing QGNNs guaranteeing efficient circuit design and scalability Quantum cryptography with GNN
@@
[…]
  • We design an overview of graph neural networks in NG IoT systems, discussing their definitions, paradigms, and key use cases. To justify the use of GNNs, we analyze the unique characteristics of NG-IoT networks and bridge them to the factor graph with the low-density parity-check (LDPC) code to show how GNNs serve as learnable generalizations of belief propagation. This connection supports exploring the GNN convergence with density evolution and training interpretability via EXIT charts. We explain the roles and benefits of node-, edge-, and graph-level tasks in wireless network problems, offering practical examples to guide their application in various NG-IoT scenarios. We compare the performance of different GNN models, including our proposed hybrid quantum GNN harnessed for power allocation in cell-free massive MIMO, demonstrating the efficiency of GNNs and the potential of quantum GNNs for future research.
  • We explore key factors influencing the efficient deployment of GNNs in real-world NG-IoT environments, including on-device implementation, multimodal data integration, cost considerations, and sustainability.
  • We present a comprehensive review of GNN applications, carefully categorized by the core technologies driving NG advances. These include massive MIMO schemes, RIS, Satellite, THz communications, MEC, URLLC, and blockchain. We also discuss the integration of GNNs with distributed systems, highlighting how they mutually enhance scalability, privacy, and collaborative intelligence in NG-IoT networks.
@@
  • Table IV offers a detailed comparison between our work and other state-of-the-art surveys in the field, emphasizing the unique contributions and advances presented in this paper.
@@
Paper Organization: The remainder of the paper is organized as follows. Section II provides an overview of GNNs, including their definitions and fundamental paradigms. We highlight the essential role of GNNs in NG-IoT networks and describe how to represent NG-IoT environments as graph structures. Various graph task types, including node-, edge-, and graph-level, are introduced, followed by a simulation study on power allocation in cell-free massive MIMO systems to compare the performance of different GNN models. In Section III, we survey GNN applications across key NG technologies, including massive MIMO, RIS, satellite communications, THz, MEC, URLLC, and blockchain, along with their integration with distributed systems. Section IV delves into adversarial GNNs, examining potential attack methods and defense strategies in GNN-based NG networks, with a focus on both homogeneous and heterogeneous graph settings. In Section V, we discuss the integration of GNNs with emerging NG technologies, including ISAC, SAGSINs, and quantum graph neural networks. Section VI will provide a set of generic design guidelines. Finally, Section VII concludes the paper. For convenience, we treat each of the eight open questions in the form of an identical structure: 1) Background; 2) State-of-the-art; 3) Challenges and future directions.
II Foundations of GNNs for NG-IoT